OpenStudy (anonymous):

Soybean meal is 16% protein, Cornmeal is 8% protein How many pounds of each should be mixed together to get 320 lbs thst is 14% protein?

OpenStudy (watchmath):

\(x+y=320\) \(.16x+.08y=.14(320)\) Try to continue from there

OpenStudy (anonymous):

These type of questions appear in one-variable algebra courses. x and y?

OpenStudy (anonymous):

I'm trying to figure out how many lbs of corn meal & soy bean there are in the mixture

OpenStudy (anonymous):

Oh, you solve one eq for x or y, then put it in the other. The x is soybean, y cornmeal.

OpenStudy (anonymous):

\[.16x+.08(320-x)=.14\times 320=44.8\] \[.16x+25.6-.08x=44.8\] \[.08x+25.6=44.8\] \[.08x=19.2\] \[x=\frac{19.2}{.08}=240\]